Given the following information, formulate an inventory management system and determine the annual order cost:

Item cost

$5.00

Order cost

$315.00

Annual holding cost (per unit per year)

       $5.00

Annual demand

26,500

A.

$3,290.74

B.

$4,568.23

C.

$4,615.40

D.

$4,543.45

E.

$3,730.89
Answer #1

Answer: - Option"B" --- $4,568.23

Explanation: - According to given data

EOQ = √2*26,500*$315/$5 = 1827.2930

Where annual order cost = (Annual usage / order size) * oder cost

= (26,500 / 1827.2930)*315 = $4,568.23